OpenStudy (anonymous):

a rectangular city park has n area of 211200 square feet. the length of the park is 660 feet. what is the width of the park?

OpenStudy (anonymous):

Area of rectangle = lxb Area= 211200 l=660 b=A/l=211200/660=320

OpenStudy (anonymous):

thanks